Mary is half Tom's age. In seven more years she will be two thirds Tom's age. When Mary is as old as Tim is now he will be 3 times as old as she is now. How old are mary and tom now?
Answer by ankor@dixie-net.com(22740)

Mary is half Tom's age.
M = .5T
:
In seven more years she will be two thirds Tom's age.
(M+7) = (2/3)(T+7)
Multiply equation by 3 to get rid of the denominator
3(M+7) = 2(T+7)
3M + 21 = 2T + 14
Substitute.5T for M:
3(.5T) + 21 = 2T + 14
1.5T - 2T = 14 - 21
-.5T = -7
T = 14 is Tom's age, Then Mary is 7 yrs old now
:
When Mary is as old as Tim is now he will be 3 times as old as she is now. How old are Mary and Tom now
:
Now Tim enters the problem, but we already know the age of Mary & Tom??